OpenStudy (anonymous):

Add the complex expressions, and simplify. (9 + 9.4i) + (-8.6 – 4i) (Points : 3) 5.8i 0.4 + 5.4i -5 5.8

OpenStudy (anonymous):

(9+9.4i) + (-8.6-4i)= (9-8.6)+i(9.4-4)= 0.4+5.4i

OpenStudy (anonymous):

9 -8.6 – 4i+ 9.4i= 0.4+5.4i

hero (hero):

\[\color {green} {\huge{\checkmark}}\]